### v1

@@ 2.1 An unstructured Bayesian model
where $N(\theta_{i}|y_{i},D_{i})$ denotes the pdf of a normal distribution with mean $y_{i}$ and variance $D_{i}$. The main appeal or the primary reason for our using the noninformative uniform prior for $\theta_{1},\cdots,\theta_{m}$ over the $m-$dimensional Euclidean space in equation (4) follows from the fact that a reasonable (in particular, HPD, highest posterior density) $1-\gamma$ credible interval for $\theta_{i}$ is identical to confidence interval $I_{i}$ for $\theta_{i}$ that was the cornerstone in the Klein et al. (2020) solution. If we denote the Cartesian product, $I_{1}\times\cdots\times I_{m}$ of $I_{i}$’s by $\Theta_{C,u}$, then it will yield a joint credible set for $\theta$ that is identical to the joint confidence set constructed by Klein et al. (2020). We will use this credible set to create one of our solutions for the rank vector $\check{r}$. While the unstructured Bayes model helps us to reproduce the frequentist confidence set of KWW, this model fails to adequately address the impact of one or more large sampling variances $D_{i}$ on the credible set of the rank vector $\check{r}$. Indeed, if $y_{i}>y_{j}$, then, irrespective of their sampling variances, $P[\theta_{i}>\theta_{j}|y]>1/2$. In other words, if $y_{1}$ is the largest of the $y$-values, then even if its sampling variance $D_{1}$ is also the largest, with a posterior probability more than 0.5, $\theta_{1}$ will exceed any other $\theta_{i}$. In particular, we noted in our applications below, that in such cases the posterior probability of the true rank $\check{r}_{1}=m$ turned out to be more than the posterior probability that $\check{r}_{i}=m$ for any $i\neq 1$. Berger and Deely (1988) and Morris and Christiansen (1996) criticized this result as unsatisfactory.
2.2 A hierarchical Bayesian model
To address the perceived weakness of the unstructured Bayes model, we augment the sampling model (3) by specifying a suitable model for $\theta_{1},\cdots,\theta_{m}$. This corresponds to a popular hierarchical model given by
@@ 2.2 A hierarchical Bayesian model
    Conditional on model parameters $\beta,A$, subpopulation means $\theta_{i}$’s are independently distributed, given by $\theta_{i}|\beta,A\stackrel{{\scriptstyle ind}}{{\sim}}N(x_{i}^{T}\beta,A),i=1,\cdots,m$,
where $\beta$ is a $p\times 1$ regression coefficient, $A$ is the model variance, and the $p\times 1$ known vector $x_{i}$ includes covariates associated with the $i$th subpopulation. The above hierarchical model has been used by many, including Efron and Morris (1975), Fay and Herriot (1979), Berger (1985), Berger and Deely (1988) and Morris and Christiansen (1996). The model (II) above provides a structure for the means $\theta_{i}$’s through the covariate $x_{i}$. If $x_{i}=1$, it corresponds to an intercept only model, which results in an exchangeable structure for the $\theta_{i}$’s. If the model error variance $A$ is small, then the covariates $x_{i}$’s introduce a lot of structure to the $\theta_{i}$’s. On the other hand, if $A$ is large, there will be little structure for the $\theta_{i}$’s. Indeed, in the limit $A\rightarrow\infty$, the model (II) does not control any structure for the $\theta_{i}$’s through $x_{i}$’s, it would be identical to the unstructured Bayesian model for $\theta_{i}$’s in Section 2.1. For known model parameters $\beta,A$, the Bayes estimators of $\theta_{i}$’s are obtained from the posterior pdf of $\theta_{i}$’s, which are independent normal distributions. The posterior pdf’s depend on the parameters $\beta,A$. Efron and Morris (1975) suggested estimate the parameters from the marginal pdf of $Y_{1},\cdots,Y_{m}$ (obtained from the joint pdf of $Y_{1},\cdots,Y_{m},\theta_{1},\cdots,\theta_{m}$ from (I) and (II) above), and use the estimated posterior pdf to make inference about $\theta_{i}$’s. The resulting solution is known as the EB solution. On the other hand, following Lindley and Smith (1972), Berger (1985), Berger and Deely (1988), and many others, we can pursue a fully Bayesian approach, also known as the HB approach, by specifying a prior pdf, typically, noninformative and improper, for $\beta,A.$ In our applications, we use the improper prior pdf
@@
where $\bar{D}$ is the mean of the $D_{i}$’s. While we use the arithmetic mean of $(D_{i}+A)$’s to define our prior pdf, Berger and Deely (1988) used the geometric mean of $(D_{i}+A)$’s to define their pdf. Aitkin and Longford (1986), Laird and Louis (1989) and Morris and Christiansen (1996) used EB or the HB estimates of $\theta_{i}$’s in place of $Y_{i}$’s for estimating the individual ranks $\check{r}_{i}$’s of the means and probability that $\theta_{i}$ is the largest of the means. The EB solutions to these problems have explicit expressions. Berger and Deely (1988) showed that HB implementation depended on evaluating several one-dimensional integrals with respect to $A$.
@@
To review the solution by Klein et al. (2020) for joint confidence set of overall ranking, we start with the confidence interval $[L_{i},U_{i}]$ for $\theta_{i}$, based on $Y_{i}$, for $i=1,\cdots,m$. Here the intervals are such that the probability $P[L_{i}\leq\theta_{i}\leq U_{i},i=1,\cdots,m]$ is as close as possible but not less than $(1-\alpha)$. For each $i\in\{1,2,\cdots,m\}$, they define the sets
@@
Using these subsets, Klein et al. (2020) in Section 3 show that the set of rank vectors
@@
is a joint confidence set for the overall rank vector $\check{r}=(\check{r}_{1},\cdots,\check{r}_{m})$ having joint coverage probability of at least $(1-\alpha)$. In the above, for a set $G\subset\{1,2,\cdots,m\}$, we use $|G|$ to denote the number of elements in $G$. Note that the joint confidence set for the vector $\check{r}$ is specified by the sets $\Lambda_{Li},\Lambda_{Oi},i=1,\cdots,m$, and in turn, these sets depend on the joint confidence sets of $\theta_{i}$’s, which themselves are determined by the estimators $Y_{i}$’s and their measures of uncertainty. In the preceding frequentist approach, the ranking is indirect through the relative positioning of individual confidence intervals. This is unlike the Bayesian approach we propose below where $\theta$ is generated from the posterior distribution and the ranking is direct based on obtained $\theta$ samples. The frequentist algorithm above relies very much on comparison of confidence intervals of components of $\theta$. The algorithm requires that a joint confidence set must be an orthotope (hyperrectangular) so that each component of $\theta$ can vary freely without being constrained by the other components. To its disadvantage, the algorithm does not allow use of more efficient joint confidence regions, for example, likelihood-based regions, since these regions are usually defined through non-linear relationship(s) among $\theta_{i}$’s.
4 Construction of credible distribution of overall ranking
For all our Bayesian models we implement our Bayesian credible solution to the overall ranking problem by a sampling-based approach. Our approach is novel, where we generate a large sample of size $S$ of $\theta_{1},\cdots,\theta_{m}$ values from their joint posterior pdf, $\pi_{M}(\theta|y)$, derived under the model $M$, where $M$ is $UB$ for unstructured, and $HB$ for hierarchical model. Generically, we denote these posterior sample values of $\theta=(\theta_{1},\cdots,\theta_{m})^{T}$ by the set $\{\theta^{(1)},\cdots,\theta^{(S)}\}$, where $\theta^{(s)}$ is an $m\times 1$ vector. For our unstructured Bayesian model in Subsection 2.1, any posterior sample $\theta^{(s)}$ is a vector of $m$ independent normal variables with mean $y_{i}$ and variance $D_{i}$, $i=1,\cdots,m$. For the HB model in Subsection 2.2, we draw these samples via Gibbs sampling (see, Gelfand and Smith (1990)), which in our applications is straightforward. In fact, for this setup it immediately follows that the full set of conditional pdf’s of $\theta_{1},\cdots,\theta_{m}$, $\beta$ and $A$, that are needed to implement Gibbs sampling, the first two distributions are based on multivariate normal distributions. To sample from the conditional pdf of $A$ we use the rejection sampling. A point estimator $\hat{\theta}$ of $\theta$ itself does not usually provide any uncertainty (for example, measure of variability) associated with it. Any reasonable confidence set (or credible set, in Bayesian approach) for $\theta$ based on $\hat{\theta}$ is an acceptable way to account for variability of $\hat{\theta}$. Usually, an appropriately constructed confidence set based on the distribution of $\hat{\theta}$ will be “tight” if the variability is small and vice versa. Based on any point estimator $\hat{\theta}$ of the vector of means $\theta$ we can get an estimator $\widehat{\check{r}}$ of the rank vector $\check{r}$ of the population means. An estimator $\widehat{\check{r}}$ is only a point estimator. Just as the estimator $\hat{\theta}$ by itself does not provide an estimator of variability, the estimator $\widehat{\check{r}}$ also does not provide any measure of variability. If $\hat{\theta}$ is accurate, so will usually be the estimator $\widehat{\check{r}}$. If $\widehat{\check{r}}$ is not accurate, it may be much different from the true rank vector $\check{r}$. Thus, it makes sense to use a better estimator of $\theta$, for example, model based EB estimator $\hat{\theta}_{EB}$ or HB estimator $\hat{\theta}_{HB}$, than the direct estimator $Y$. But any point estimator of rank $\check{r}$ based on a model based estimator of $\theta$ still does not account for uncertainty. Estimators of individual ranks based on some model based estimators of $\theta$ have been considered by Aitkin and Longford (1986), Berger and Deely (1988), Laird and Louis (1989) or Morris and Christiansen (1996). However, such estimators do not account for uncertainty of the joint overall ranking. Klein et al. (2020) noted that assessment of the uncertainty in the estimated overall ranking would involve all the populations simultaneously and their relative standing to each other. They are the first to present overall ranking of population means that accounts for uncertainty of the released rankings of population means. They adopted a frequentist approach and achieved this by deriving a joint confidence set for rank vector $\check{r}$ from an appropriate joint confidence set for the mean vector $\theta$. In particular, if the confidence set for reasonable confidence coefficient is “tight”, the resulting inference for $\check{r}$ will be quite accurate. Klein et al. (2020) first created a joint confidence set of the form $I_{1}\times\cdots\times I_{m}$ for the mean vector $\theta$, where $I_{i}$ is a suitable confidence interval for $\theta_{i}$ so that the joint confidence level of the set $I_{1}\times\cdots\times I_{m}$ is $(1-\alpha)$. They chose $I_{i}=(Y_{i}-z_{1-\gamma/2}\sqrt{D_{i}},Y_{i}-z_{\gamma/2}\sqrt{D_{i}})\equiv(L_{i},U_{i})$ where $\gamma$ is so chosen that a joint confidence level $(1-\alpha)$ is achieved. From this joint confidence set, in Section 3 of their paper, they constructed joint confidence set of the overall ranking vector $\check{r}$. We reviewed this solution in Section 3. In our Bayesian approach to joint credible set for overall ranking, as it was done by KWW in a frequentist approach, we start with a suitable joint credible set for $\theta$. However, our approach to creating the credible set for overall ranking vector is much different and, we think, our solution is more informative than the KWW solution. To explain our ideas in simple terms, we consider $\theta_{1}$, which is a scalar. Based on a large number of sampled values $\theta_{1}^{(1)},\cdots,\theta_{1}^{(S)}$ from the posterior distribution of $\theta_{1}$, to empirically create a reasonable $1-\delta$ credible interval for $\theta_{1}$, we can use the $(\delta/2)$th and the $(1-\delta/2)$th quantiles of the sampled values. Intuitively, all the $\theta_{1}^{(s)}$ values which are inside this empirical credible interval may be considered “plausible values” of $\theta_{1}$. By extending this idea analogously to the whole vector $\theta$, first we draw a sample $\theta^{(1)},\cdots,\theta^{(S)}$ from the joint posterior distribution $\pi_{M}(\theta|y)$ of $\theta$. Suppose ${\cal F}_{M}$ denotes this sample. Next we construct empirically a suitable joint $(1-\alpha)$ credible set, ${\cal F}_{M,\alpha,y}$, for $\theta$ based on the samples ${\cal F}_{M}$. We present construction of two intuitively appropriate credible sets below. It is reasonable to consider that the elements of ${\cal F}_{M,\alpha,y}$ are plausible values of $\theta$. We will use these plausible values to create a credible distribution (explained below) for the overall rank vector $\check{r}$.
4.1 Finding an empirical joint credible set for $\theta$
@@
4.1.1 A Cartesian credible set for $\theta$
@@
This approach closely resembles the confidence set constructed by Klein et al. (2020). From the vectors in ${\cal F}_{M}$, for $i=1,\cdots,m$, we take the $i$th component of each vector to create a sample ${\cal F}_{i}$ for $\theta_{i}$ from its marginal posterior distribution. For some suitable $\kappa\in(0,1)$, we determine the $(\kappa/2)$th and $(1-\kappa/2)$th quantiles, $a_{i}$ and $b_{i}$, respectively, of the set ${\cal F}_{i}$. Let ${\cal S}_{i}=\{s:a_{i}\leq\theta_{i}^{(s)}\leq b_{i},s=1,2,\cdots S\}$. Define the set ${\cal S}_{J}=\cap_{i=1}^{m}{\cal S}_{i}$. Let $K_{J}=|{\cal S}_{J}|$. For any $\alpha\in(0,1)$, we will determine $\kappa$ in such a way that $K_{J}$ is the integer closest to $S(1-\alpha)$. For this choice of $\kappa$, the set $\{\theta^{(s)},s\in{\cal S}_{J}\}$ is an empirical $(1-\alpha)$ joint credible set of $\theta$. For a given $\alpha$, an initial choice of $\kappa$ is $1-(1-\alpha)^{1/m};$ however, we need to iteratively modify $\kappa$ to achieve the approximate relation $K_{J}\approx S(1-\alpha)$ as closely as possible. Finally, corresponding to the selected set $\{\theta^{(s)},s\in{\cal S}_{J}\}$, we obtain an approximate $(1-\alpha)$ joint credible distribution for the rank vector $\check{r}$. In Remark 8 of their paper, Klein et al. (2020) suggested constructing a Cartesian credible set for $\theta$ and using this credible set to create a credible set for the overall ranking by applying the main result of their paper. We note that the joint credible set for $\theta$ constructed this way by these authors for the Bayesian model in Section 2.1 will be identical to their frequentist confidence set, but this time the result will be interpreted conditional on the data instead of conditional on the $\theta$ for the frequentist case. However, we show by simulations and a real application to median income estimation (in Section 9) that our Bayesian solution to ranking produces a smaller error in estimation measured by expected posterior absolute deviation from the “true ranking” or a “gold standard” for it.
4.1.2 An elliptical credible set for $\theta$

### v2

@@ 2.1 An unstructured Bayesian (UB) model
where $N(\theta_{i}|y_{i},D_{i})$ denotes the pdf of a normal distribution with mean $y_{i}$ and variance $D_{i}$. The main reason for the noninformative uniform prior for $\theta_{1},\cdots,\theta_{m}$ over the $m$-dimensional Euclidean space in equation (4) follows from the fact that a reasonable (in particular, HPD, highest posterior density) $1-\gamma$ credible interval for $\theta_{i}$ is identical to confidence interval $I_{i}$ for $\theta_{i}$ that was the cornerstone in the Klein et al. (2020) solution. If we denote the Cartesian product, $I_{1}\times\cdots\times I_{m}$ of $I_{i}$’s by $\Theta_{C,u}$, then it will yield a joint credible set for $\theta$ that is identical to the joint confidence set constructed by Klein et al. (2020).
We will use this credible set to create one set of credible distributions solutions for the rank vector $\check{r}$. While the unstructured Bayes model helps us to reproduce the frequentist confidence set of KWW, this model fails to adequately address the impact of one or more large sampling variances $D_{i}$ on the credible set of the rank vector $\check{r}$. Indeed, if $y_{i}>y_{j}$, then, irrespective of their sampling variances, $P[\theta_{i}>\theta_{j}|y]>1/2$. In other words, if $y_{1}$ is the largest of the $y$-values, then even if its sampling variance $D_{1}$ is the largest and excessively large, with a posterior probability more than 0.5, $\theta_{1}$ will exceed any other $\theta_{i}$. In particular, we noted in our applications below, that in such cases the posterior probability of $\check{r}_{1}=m$ turned out to be more than the probability that $\check{r}_{i}=m$ for any $i\neq 1$. Berger and Deely (1988) and Morris and Christiansen (1996) justifiably criticized this result as unsatisfactory.
2.2 A hierarchical Bayesian (HB ) model
To address the perceived weakness of the UB model, we augment model (3) by specifying a suitable model for $\theta_{1},\cdots,\theta_{m}$. This corresponds to a popular hierarchical model:
@@ 2.2 A hierarchical Bayesian (HB ) model
    $\theta_{i}|\beta,A\stackrel{{\scriptstyle ind}}{{\sim}}N(x_{i}^{T}\beta,A),i=1,\cdots,m$,
where $\beta$, a $p\times 1$ regression coefficient, and $A$ the model variance, are the model parameters. The $p\times 1$ known vector $x_{i}$ includes covariates associated with the $i$th subpopulation. This model was used by many, including Efron and Morris (1975), Berger (1985), Berger and Deely (1988) and Morris and Christiansen (1996). This model was first used by Fay and Herriot (1979) for EB in small area estimation (SAE). In the SAE literature this is the celebrated Fay-Herriot model that is extensively used for EBLUP and EB approaches to SAE, e.g., p. 77 of Rao and Molina (2015).
The step (II) above, links $\theta_{i}$ to $x_{i}$, provides a structure for the means $\theta_{i}$’s through the covariate $x_{i}$. It is known as the linking model in SAE, see again Rao and Molina (2015). If $x_{i}=1$, it corresponds to an intercept only model, which results in an exchangeable structure for the $\theta_{i}$’s. If the model error variance $A$ is small, then the covariates $x_{i}$’s introduce a lot of structure to the $\theta_{i}$’s. On the other hand, if $A$ is large, there will be little structure for the $\theta_{i}$’s. Indeed, if $A\rightarrow\infty$, the model (II) does not control any structure for the $\theta_{i}$’s through $x_{i}$’s, it would be identical to the UB model for $\theta_{i}$’s in Section 2.1.
For known model parameters $\beta,A$, the Bayes estimators of $\theta_{i}$’s are obtained from the posterior pdf of $\theta_{i}$’s, which are independent normal distributions. The posterior pdf’s depend on $\beta$ and $A$. Efron and Morris (1975) suggested estimate the parameters from the marginal pdf of $Y_{1},\cdots,Y_{m}$ (obtained from the joint pdf of $Y_{1},\cdots,Y_{m},\theta_{1},\cdots,\theta_{m}$ from (I) and (II) above), and use the estimated (based on estimates of model parameters) posterior pdf to make inference about $\theta_{i}$’s. The resulting solution is known as the EB solution.
On the other hand, Berger (1985), Berger and Deely (1988), and many others, pursued a fully Bayesian approach, the HB approach, by specifying a prior pdf, typically, noninformative and improper, for $\beta,A.$ In our applications, we use the improper prior pdf
@@
where $\bar{D}=\sum_{i=1}^{m}D_{i}/m$. While we use $\bar{D}+A$ to define our prior pdf, Berger and Deely (1988) used the geometric mean of $(D_{i}+A)$’s to define their pdf.
Aitkin and Longford (1986), Laird and Louis (1989) and Morris and Christiansen (1996) used EB or HB estimates of $\theta_{i}$’s in place of $Y_{i}$’s for estimating the individual ranks $\check{r}_{i}$’s and probability that $\theta_{i}$ is the largest of the means. The EB solutions to these problems have explicit expressions. Berger and Deely (1988) showed that HB solution depended on evaluating several one-dimensional integrals with respect to $A$. The hierarchical model in (I) and (II), complemented by a prior pdf on $\beta,A$, as in equation (6), is known as the Fay-Herriot HB model in the SAE literature.
@@
To review Klein et al. (2020) solution for joint confidence set of overall ranking, we start with the confidence interval $[L_{i},U_{i}]$ for $\theta_{i}$, based on $Y_{i}$, for $i=1,\cdots,m$. These intervals are such that the (frequentist) probability $P\{\cap_{i=1}^{m}[L_{i}\leq\theta_{i}\leq U_{i}]\}$ is no less than $(1-\alpha)$. For each $i\in\{1,2,\cdots,m\}={\cal M}$, they define various subsets of ${\cal M}$, given below,
@@
Using these subsets, in Section 3 Klein et al. (2020) show that the set of rank vectors
@@
is a joint confidence set for the overall rank vector $\check{r}=(\check{r}_{1},\cdots,\check{r}_{m})$ having a joint coverage probability at least $(1-\alpha)$. For a set $G\subset{\cal M}$, the number of elements in $G$ is $|G|$. The joint confidence set for the vector $\check{r}$ is specified by the sets $\Lambda_{Li},\Lambda_{Oi},i=1,\cdots,m$, which depend on the joint confidence sets of $\theta_{i}$’s, that are determined by the estimators $Y_{i}$’s and their measures of uncertainty.
@@
In the frequentist approach above, the ranking is indirect through the relative positioning of individual confidence intervals. This is unlike the credible distributions (CDs) under the Bayesian approach we shortly propose where samples of $\theta$ are generated from the posterior distribution and the ranking is direct based on $\theta$ samples. The frequentist algorithm above relies very much on comparison of confidence intervals of components of $\theta$. It requires that a joint confidence set must be an orthotope (hyperrectangular) so that each component of $\theta$ can vary freely without being constrained by the other components. To its disadvantage, the algorithm does not allow use of more efficient joint confidence regions, for example, likelihood-based regions, since ranges of $\theta_{i}$’s in these regions are usually interdependent.
4 Description of empirical credible sets
For our Bayesian models we create credible distributions by a sampling-based approach. In this approach we generate a large sample of size $S$ of $\theta_{1},\cdots,\theta_{m}$ values from their joint posterior pdf, $\pi_{M}(\theta|y)$, derived under the model $M$, where $M$ is $UB$ for unstructured, and $HB$ for hierarchical model. Generically, we denote these posterior sample values of $\theta=(\theta_{1},\cdots,\theta_{m})^{T}$ by the set $\{\theta^{(1)},\cdots,\theta^{(S)}\}$, where $\theta^{(s)}$ is an $m\times 1$ vector.
For the UB model in Subsection 2.1, any posterior sample $\theta^{(s)}$ is a vector of $m$ independent normal variables with mean $y_{i}$ and variance $D_{i}$, $i=1,\cdots,m$. For the HB model in Subsection 2.2, we draw these samples via MCMC, which in our applications is straightforward. See Subsection S.1 in the Supplement for details.
A point estimator $\hat{\theta}$ by itself does not usually provide any uncertainty (for example, measure of variability) associated with it. Any reasonable confidence set (or credible set, in Bayesian approach) for $\theta$ is an acceptable way to account for variability of $\hat{\theta}$. Usually, an appropriately constructed confidence set based on the distribution of $\hat{\theta}$ will be “tight” if the variability of $\hat{\theta}$ is small and vice versa.
Based on any point estimator $\hat{\theta}$ we can get a point estimator $\widehat{\check{r}}$ of the rank vector $\check{r}$. Just as the estimator $\hat{\theta}$ by itself does not provide an estimator of variability, the estimator $\widehat{\check{r}}$ also does not provide any measure of variability. If $\hat{\theta}$ is accurate, so will usually be the estimator $\widehat{\check{r}}$. If $\widehat{\check{r}}$ is not accurate, it may be much different from the true rank vector $\check{r}$. Thus, it makes sense to use a better estimator of $\theta$, for example, model-based EB estimator $\hat{\theta}_{EB}$ or HB estimator $\hat{\theta}_{HB}$, than the direct estimator $Y$. But any point estimator of rank $\check{r}$ based on a model-based estimator of $\theta$ still does not account for uncertainty.
Estimators of individual ranks based on model-based estimators of $\theta$ have been considered by Aitkin and Longford (1986), Berger and Deely (1988), Laird and Louis (1989), Goldstein and Spiegelhalter (1996) or Morris and Christiansen (1996). Like ours, the approach to this problem by Berger and Deely (1988) and Goldstein and Spiegelhalter (1996) is fully Bayesian. The last group of authors considered a Bayesian approach to comparison of several health or educational institutions. They developed for their applications 95% credible intervals for individual ranks. However, none of these estimation methods considers joint overall ranking. Klein et al. (2020) noted that assessment of the uncertainty in the estimated overall ranking would involve all the populations simultaneously and their relative standing to each other.
For our Bayesian solutions we follow KWW only to construct a suitable joint credible set for $\theta$. Subsequently, our approach to creating set estimate for overall ranking vector is much different and, evidence shows, our solution is more informative. To explain our ideas in simple terms, we consider $\theta_{1}$, which is a scalar. Based on a large number of sampled values $\theta_{1}^{(1)},\cdots,\theta_{1}^{(S)}$ from the posterior distribution of $\theta_{1}$, their $(\delta/2)$th and $(1-\delta/2)$th quantiles reasonably provide an empirical $1-\delta$ credible interval for $\theta_{1}$. Intuitively, all the $\theta_{1}^{(s)}$ values which are inside this empirical credible interval may be considered “plausible values” of $\theta_{1}$. By extending this idea analogously to the whole vector $\theta$, first we draw ${\cal F}_{M}=\{\theta^{(1)},\cdots,\theta^{(S)}\}$, a sample from the joint posterior distribution $\pi_{M}(\theta|y)$ of $\theta$. Next we construct empirically a suitable joint $(1-\alpha)$ credible set, ${\cal F}_{M,\alpha,y}$, for $\theta$ based on the samples ${\cal F}_{M}$. We present construction of two intuitively appropriate credible sets below. It is reasonable to consider that the elements of ${\cal F}_{M,\alpha,y}$ are plausible values of $\theta$. We will use these plausible values to create credible distributions (explained below) for the ranks $\check{r}_{1},\cdots,\check{r}_{m}$.
4.1 Constructing ${\cal F}_{M,\alpha,y}$: an empirical joint credible set for $\theta$
@@
4.1.1 A Cartesian empirical credible set for $\theta$
@@
This approach closely follows only the confidence set construction by Klein et al. (2020) for $\theta$; our credible solution for $\check{r}$ is different. From the vectors in ${\cal F}_{M}$, for $i=1,\cdots,m$, we take the $i$th component of each vector to create a sample ${\cal F}_{i}$ for $\theta_{i}$ from its marginal posterior distribution. For some suitable $\kappa\in(0,1)$, we determine the $(\kappa/2)$th and $(1-\kappa/2)$th quantiles, $a_{i}$ and $b_{i}$, respectively, of the set ${\cal F}_{i}$. Let ${\cal S}_{i}=\{s:a_{i}\leq\theta_{i}^{(s)}\leq b_{i},s=1,2,\cdots S\}$. Define the set ${\cal S}_{J}=\cap_{i=1}^{m}{\cal S}_{i}$. Let $K_{J}=|{\cal S}_{J}|$. For any $\alpha\in(0,1)$, we will determine $\kappa$ in such a way that $K_{J}$ is the integer closest to $S(1-\alpha)$. For this choice of $\kappa$, the set $\{\theta^{(s)},s\in{\cal S}_{J}\}$ is an empirical $(1-\alpha)$ joint credible set of $\theta$. For a given $\alpha$, an initial choice of $\kappa$ is $1-(1-\alpha)^{1/m};$ however, we need to iteratively modify $\kappa$ to satisfy the approximate relation $K_{J}\approx S(1-\alpha)$ as closely as possible. Finally, corresponding to the selected set $\{\theta^{(s)},s\in{\cal S}_{J}\}$, we obtain an approximate $(1-\alpha)$ joint credible distribution for the rank vector $\check{r}$.
In Remark 8 of their paper, Klein et al. (2020) suggested constructing a Cartesian credible set for $\theta$ and using this set to create a credible set for the overall ranking by applying the main result of their paper. We note that the joint credible set for $\theta$ constructed this way by these authors for the Bayesian model in Section 2.1 will be identical to their frequentist confidence set, but this time the result will be interpreted conditional on the data instead of conditional on the $\theta$ as is done in the frequentist case. However, we show by simulations and a real application to median income estimation (in Section 9) that our Bayesian solution to ranking produces a smaller error in estimation measured by expected posterior absolute deviation from the “true ranking,” or some “gold standard” for it.
4.1.2 An elliptical empirical credible set for $\theta$